1. Under any normal distribution of scores, what percentage of the total area falls

A. between the mean (μ) and a score value that lies one standard deviation (1σ) above the mean?

B. between a score value that lies one standard deviation below the mean and a score value that lies one standard deviation above the mean?

C. between the mean and a score value that lies +2σ above the mean?

D. between a score value that lies −2σ below the mean and a score value that lies +2σ above the mean?
Answer #1

The empirical rule states that for a normal distribution, nearly all of the data will fall within three standard deviations of the mean. The empirical rule can be broken down into three parts:

  • 68% of data falls within the first standard deviation from the mean.
  • 95% fall within two standard deviations.
  • 99.7% fall within three standard deviations.

The rule is also called the 68-95-99 7 Rule or the Three Sigma Rule.

one standard deviation 68% of data — - 95% of data 99.7% of data

A )

The percentage of the total area falls between the mean (μ) and a score value that lies one standard deviation (1σ) above the mean = 68 / 2 = 34%

B)

The percentage of the total area falls  between a score value that lies one standard deviation below the mean and a score value that lies one standard deviation above the mean = 68%

C)

The percentage of the total area falls between the mean and a score value that lies +2σ above the mean = 95 / 2 = 47.5%

D)

The percentage of the total area falls between a score value that lies −2σ below the mean and a score value that lies +2σ above the mean = 95%